# Homogeneous braids are visually prime

### v1

@@ 4.4. Arborescent fibered links are prime
[…]
@@
The vertices of $T$ are chosen to be open books $(A,\tau^{\pm})$, where $A$ is an annulus and $\tau^{\pm}$ is a positive/negative Dehn twist, with sign chosen according to the label. The edges are chosen according to the edges of $G$. The choice of labeling of the edges is where all the subtlety lies. The fact is that the plane tree $G$ induces a circular order on the edges at a vertex of $G$ (and hence of $T$). We choose the plumbing regions $P_{1},\ldots,P_{n}$ in an annulus $A$ that is the surface of a vertex $v=(A,\tau^{\pm})$ of $T$ the we take as the labels of the edges $e_{1}$ to $e_{n}$ that contain $v$ to be $P_{1},\ldots,P_{n}$ in any order such that the circular order induced by the orientation of $A$ equals the circular order induced by $G$. This process yields an open book well-defined up to equivalence, as the following argument shows. Any other choice of order of the $P_{i}$ that also respects the circular order described above, amounts to changing the plumbing regions by applying an orientation preserving diffeomorphism of $A$, hence resulting in a tree of open books for which the corresponding open book is equivalent to that of $T$.
@@
Note that the without a choice of circular order at each vertex, e.g. if $G$ was taken to be a $\pm$–labeled tree without preferred embedding into the plane, different choices of plane embeddings do actually lead to non-isotopic links in $S^{3}$ [Ger06].
@@ 5.1. Preliminaries on decompositions and braids
For a link $L$, we call an embedded $2$–sphere $S\subset S^{3}$ a decomposition sphere, if $S$ intersects $L$ transversely as follows: writing $S^{3}=B_{1}\cup B_{2}$ with $B_{i}\subset S^{3}$ compact balls with boundary $S$, for both $i=1$ and $i=2$, the closed arc component $a$ of $L\cap B_{i}$ is not boundary parallel in $B_{i}\setminus(L\setminus a)$. In other words, for both $i=1$ and $i=2$, $B_{i}\cap L$ is not isotopic to the result of split sum of a trivial $1$–tangle with a link. We call a link prime, if it does not admit a decomposition sphere. We note that in this definition, the unknot is a prime link, and a split link is prime if and only if its split components are prime.
@@
Let $W$ be a braid word on $n$ strands, and let $11\leq i\leq j\leq n-$ 1. Then we denote by $W_{ij}$ the braid word on $j-i+2$ strands obtained from $W$ by deleting all $\sigma_{k}^{\pm 1}$ with $k<i$ or $k>j$, and shifting the index down by $i-1$, i.e. replacing each $\sigma_{k}^{\pm 1}$ with $i\leq k\leq j$ by $\sigma_{k-i+1}^{\pm 1}$.
The word $W$ is called split if for some $i\in\{1,\ldots,n-1\}$, there are no $\sigma_{i}^{\pm 1}$ in $W$. If $W$ contains $\sigma_{n-1}^{\pm 1}$ precisely once, then one may produce the $(n-1)$–stranded word $W_{1,n-2}$ by deleting $\sigma_{n-1}^{\pm 1}$. Similarly, if $W$ contains $\sigma_{1}^{\pm 1}$ precisely once, one may produce $W_{2,n-1}$ by deleting $\sigma_{1}^{\pm 1}$ and shifting the index down by $1$. This is called (Markov) destabilization, and its inverse (Markov) stabilization. Stabilization and destabilization leave the closure invariant.
@@ 5.2. Decomposition circles of (homogeneous) braids
To analyze decomposition circles of diagrams $D(W)$, we need to count ‘how often $W$ alternates between $\sigma_{i-1}^{\pm 1}$ and $\sigma_{i}^{\pm 1}$ generators’. More precisely, for a braid word $W$ on $n$ strands and $12\leq i\leq n-$ 1, let the $i$–th seesaw number $g_{i}(W)$ be defined as follows. The word $W_{i-1,i}$ can be uniquely written as $V_{1}\cdots V_{g_{i}(W)}$ for some $g_{i}(W)\geq 0$, such that $V_{\ell}$ is a non-empty braid word consisting only of $\sigma_{i-1}^{\pm 1}$ if $\ell+\epsilon$ is even, and only of $\sigma_{i}^{\pm 1}$ if $\ell+\epsilon$ is odd, for some fixed $\epsilon\in\{0,1\}$.
@@
(i). Let $C\subset S^{2}$ be a circle transverse to $D(W)$. After an isotopy pushing $C$ away from the crossings of $D(W)$, all intersection points $C\cap D(W)$ are on one of the strands. It follows that $C$ intersects each strand an even number of times. Thus if $C$ has two intersections with $D(W)$, then $C$ intersects only one strand (let us say the $i$–th one), and does so twice. One checks that this is possible if and only if $i=1$, $i=n$, or $12\leq i\leq n-$ 1 and $g_{i}(W)\leq 3$. Moreover, $C$ is a decomposition circle if and only if $12\leq i\leq n-$ 1 and $g_{i}(W)\geq 2$.
(ii). Denote by $m_{-}$ and $m_{+}$ the number of times $W$ has been destabilized at the lower and upper end, respectively, to obtain $W^{\prime}$. Assume towards a contradiction that $g_{i}(W^{\prime})\leq 3$ for some $i$. Since $W^{\prime}$ is non-split, $g_{i}(W^{\prime})\geq 2$. Hence, as proven above, there is a decomposition circle $C^{\prime}$ of $D(W^{\prime})$ that intersects the $i$–th strand of $W^{\prime}$ twice. This gives rise to a decomposition circle $C$ of $D(W)$ that intersects the $(i+m_{-})$–th strand of $W$ twice. This circle decomposes the link $\operatorname{cl}(W)$ as a connected sum of $L_{-}\coloneqq\operatorname{cl}(W_{1,i+m_{-}-1})$ and $L_{+}\coloneqq\operatorname{cl}(W_{i+m_{-},n})$. Since $W^{\prime}$ is non-split and cannot be destabilized, $W_{1,i+m_{-}-1}$ contains at least two copies of $\sigma_{m_{-}+1}^{\pm 1}$. Similarly, $\operatorname{cl}(W_{i+m_{-},n})$ contains at least two copies of $\sigma_{n-m_{+}}^{\pm 1}$. Using that homogeneous braids are visually trivial and visually split, it follows that $L_{+}$ and $L_{-}$ are non-trivial non-split links, and thus $C$ is an honest decomposition circle of $W$. ∎
@@ 5.3. Seifert surfaces of braids as Murasugi sums
Let us now describe the Seifert surface $\Sigma(W)$ for $\operatorname{cl}(\beta)$, obtained by applying Seifert’s algorithm to $D(W)$. The surface $\Sigma(W)$ consists of disks $D_{1},\dots,D_{n}$, such that $D_{i}$ has radius $n-i+1$, center $(0,0,i)$, is parallel to the $xy$–plane, and its boundary is the $i$–th strand of $W$; as well as one twisted band for each crossing for each crossing of $W$. We call $D_{1}$ the bottom disk and $D_{n}$ the top disk. The following lemma gives a precise version of the well-known decomposition of $\Sigma(W)$ as Murasugi sum along the disks $D_{i}$. Examples can be seen in Figure 2, Figure 3, and Figure 4 in Section 2.
@@
Let us now explicitly define $P$. Recall from the definition of $g_{i}(W)\eqqcolon g$ that $W_{i-1,1}=V_{1}\cdots V_{g}$, where the words $V_{\ell}$ consist alternatingly only of $\sigma_{i-1}^{\pm 1}$, or only of $\sigma_{i}^{\pm 1}$. Now, let us pick points $p_{1},\ldots,p_{g}\in\partial\Sigma(W)$ that lie in this order on the circle $\partial D_{i}$, such that $p_{j}$ is between the spots where the twisted band corresponding to the last letter in $V_{j}$ and the twisted band corresponding to the first letter in $V_{j+1}$ are affixed to $D_{i}$ (where we see indexes modulo $g$ so that $V_{g+1}=V_{1}$). Let the polygon $P$ now be given a subset of the disk $D_{i}$, with vertices $p_{1},\ldots,p_{g-1}$ (and $p_{g}$ is $g$ is even), and edges given as chords of $D_{i}$. One checks that after an isotopy, the sides of $P$ indeed belong alternatingly to the boundary of the lower and the upper Murasugi summand of $\Sigma(W)$, thus proving (iii).
Finally, to show (iv), let us assume $g\geq 4$. Let $1\leq j\leq g$ and $j<g$ if $g$ is odd. Let us treat indexes modulo $g$ if $g$ is even, and modulo $g-1$ if $g$ is odd. Pick two twisted band connecting $D_{i}$ and $D_{i\pm 1}$: one of them affixed to $D_{i}$ between $p_{j}$ and $p_{j+1}$, the other between $p_{j+2}$ and $p_{j+3}$. Pick a simple closed curve $\gamma_{j}$ in $\Sigma(W)$ that consists of the core curves of those two twisted bands, and a chord in each of $D_{i}$ and $D_{i\pm 1}$. Observe that $\gamma_{j}$ is either contained in the lower or in the upper Murasugi summand of $\Sigma(W)$, depending on the sign in $D_{i\pm 1}$. Moreover, $\gamma_{j}$ intersects the side of $P$ between $p_{j}$ and $p_{j+1}$ exactly once. That is sufficient to prove that this side is essential. Overall, the sides of $P$ are alternatingly essential in the two Murasugi summands of $\Sigma(W)$; hence, the Murasugi sum is essential by definition. ∎
@@
Then $W$ satisfies the hypothesis of Lemma 5.1 (ii). So, for $W^{\prime}$ obtained from $W$ by destabilizing as often as possible, we have $g_{i}(W^{\prime})\geq 4$ for all $i$. Note that $\Sigma(W)$ and $\Sigma(W^{\prime})$ are isotopic. Hence it suffices to prove that $\Sigma(W^{\prime})$ is prime. Also note that $W^{\prime}$ is itself homogeneous and non-split. Let $n$ be the number of strands of $W^{\prime}$.
An iterative application of Lemma 5.2 decomposes $\Sigma(W^{\prime})$ as an $(n-1)$–fold Murasugi sum with summands $\Sigma(W^{\prime}_{i,i})$. Because $W^{\prime}$ is homogeneous, every $W^{\prime}_{i,i}$ consists either of $k$ copies of $\sigma_{1}$, or of $k$ copies of $\sigma_{1}^{-1}$, with $k\geq 2$. So $\Sigma(W^{\prime}_{i,i})$ is the fiber surface of the $T(2,\pm k)$–torus link.
This Murasugi sum decomposition of $\Sigma(W^{\prime})$ is encoded by a tree of open books $T$ as defined in Section 3. Namely, the vertices of $T$ are the open books of the $\Sigma(W^{\prime}_{i,i})$; and there is an edge precisely between $\Sigma(W^{\prime}_{i-1,i-1})$ and $\Sigma(W^{\prime}_{i,i})$ for $12\leq i\leq n-$ 1. Note that the tree $T$ simply has the shape of a line. The edge between $\Sigma(W^{\prime}_{i-1,i-1})$ and $\Sigma(W^{\prime}_{i,i})$ is labeled by the summing region $P_{i}$ given by Lemma 5.2. Let us check that $T$ satisfies the hypothesis of a tree of open books. Firstly, the summing regions are disjoint, because $P_{i}$ is embedded in the $i$–th disk $D_{i}$ of $\Sigma(W^{\prime})$. Secondly, since $g_{i}(W^{\prime})\geq 4$ for all $i$, it follows from Lemma 5.2 (iv) that $P_{i}$ is essential. So $T$ is indeed a tree of open books. The theorem now follows from Theorem 4.5. ∎

### v2

@@ 4.4. Arborescent fibered links are prime
As an application of Theorem 4.5 and an illustration of examples of trees of open books, we discuss the primeness of arborescent fibered links. We do so without recalling the construction of arborescent links in detail, as we see this as an illustration of the technique rather than a main result of the text. In particular, since primeness for most arborescent links can be derived from most of them being hyperbolic, as discussed in [FG09], we do not claim originality but rather a different perspective. For a general reference for the construction of arborescent links (certain links associated with a plane tree with integer labeled vertices) we refer to [BS16, Ch. 12]. Fibered arborescent links are the links $L_{G}$ in $S^{3}$ (well-defined up to ambient isotopy) associated to a plane tree $G$ with $\pm$–labeled vertices, i.e. a choice of embedding of a finite tree with vertices labeled by elements in $\{+,-\}$. The link $L_{G}$ can be described by placing a Hopf band in the plane at each vertex (with the sign of the Hopf band being dictated by the sign of the corresponding vertex) and plumbing them together whenever they share an edge; compare with [Baa05] where this construction is explained, and consult [BS16, Ch. 12] for a detailed account. In the language of Murasugi sums, arborescent links are those arising from plumbing together oriented unknotted annuli (understood as open books for $S^{3}$). They are fibered if each of the involved annuli is fibered (since Murasugi sums have fibered boundary if and only if the pieces have fibered boundary [Gab83, Theorem 4]); hence, fibered arborescent links are those that arise if each annulus is a Hopf band.
@@
The vertices of $T$ are chosen to be open books $(A,\tau^{\pm})$, where $A$ is an annulus and $\tau^{\pm}$ is a positive/negative Dehn twist, with sign chosen according to the label. The edges are chosen according to the edges of $G$. The choice of labeling of the edges is where all the subtlety lies. The fact is that the plane tree $G$ induces a circular order on the edges at a vertex of $G$ (and hence of $T$). We choose the plumbing regions $P_{1},\ldots,P_{n}$ in an annulus $A$ that is the surface of a vertex $v=(A,\tau^{\pm})$ of $T$ and we take as the labels of the edges $e_{1}$ to $e_{n}$ that contain $v$ to be $P_{1},\ldots,P_{n}$ in any order such that the circular order induced by the orientation of $A$ equals the circular order induced by $G$. This process yields an open book well-defined up to equivalence, as the following argument shows. Any other choice of order of the $P_{i}$ that also respects the circular order described above, amounts to changing the plumbing regions by applying an orientation preserving diffeomorphism of $A$, hence resulting in a tree of open books for which the corresponding open book is equivalent to that of $T$.
@@
Note that without a choice of circular order at each vertex, e.g. if $G$ was taken to be a $\pm$–labeled tree without preferred embedding into the plane, different choices of plane embeddings do actually lead to non-isotopic links in $S^{3}$ [Ger06].
@@ 5. Homogeneous braids
In general, a link diagram that arises as the closure of a braid diagram representing a non-prime link will not readily allow the observation that the link is not prime. While it is known that the necessary isotopy can be chosen to be well adapted to the braid axis—namely a composition of applications of the braid relations and conjugations and so-called exchange moves suffice [BM90, The composite braid theorem]—as far as the authors know there are no a priori bounds on how many applications of braid relations, conjugations and exchange moves are needed.
In contrast, Theorem 1.1 states that if a diagram of a link arises as the closure of a braid diagram that is homogeneous, then if the link represented by the diagram is not prime, this can be seen in the diagram by means of a decomposition circle. In this section, we first establish the notation and definitions relevant to Theorem 1.1, and then prove it.
@@ 5.1. Preliminaries on decompositions and braids
For a link $L$, we call an embedded 22–sphere $S\subset S^{3}$ a decomposition sphere, if $S$ intersects $L$ transversely as follows: writing $S^{3}=B_{1}\cup B_{2}$ with $B_{i}\subset S^{3}$ compact balls with boundary $S$, for both $i=1$ and $i=2$, the closed arc component $a$ of $L\cap B_{i}$ is not boundary parallel in $B_{i}\setminus(L\setminus a)$. In other words, for both $i=1$ and $i=2$, $B_{i}\cap L$ is not isotopic to the result of split sum of a trivial 11–tangle with a link. We call a link prime, if it does not admit a decomposition sphere. We note that in this definition, the unknot is a prime link, and a split link is prime if and only if its split components are prime.
@@
Let $W$ be a braid word on $n$ strands, and let $1\leq i\leq j\leq n-1$. Then we denote by $W_{ij}$ the braid word on ji+2j-i+2 strands obtained from $W$ by deleting all $\sigma_{k}^{\pm 1}$ with $k<i$ or $k>j$, and shifting the index down by i1i-1, i.e. replacing each $\sigma_{k}^{\pm 1}$ with $i\leq k\leq j$ by $\sigma_{k-i+1}^{\pm 1}$.
The word $W$ is called split if for some $i\in\{1,\ldots,n-1\}$, there are no $\sigma_{i}^{\pm 1}$ in $W$. If $W$ contains $\sigma_{n-1}^{\pm 1}$ precisely once, then one may produce the (n1)(n-1)–stranded word $W_{1,n-2}$ by deleting $\sigma_{n-1}^{\pm 1}$. Similarly, if $W$ contains $\sigma_{1}^{\pm 1}$ precisely once, one may produce $W_{2,n-1}$ by deleting $\sigma_{1}^{\pm 1}$ and shifting the index down by $1$. This is called (Markov) destabilization, and its inverse (Markov) stabilization. Stabilization and destabilization leave the closure invariant.
@@ 5.2. Decomposition circles of (homogeneous) braids
To analyze decomposition circles of diagrams $D(W)$, we need to count ‘how often $W$ alternates between $\sigma_{i-1}^{\pm 1}$ and $\sigma_{i}^{\pm 1}$ generators’. More precisely, for a braid word $W$ on $n$ strands and $2\leq i\leq n-1$, let the ii–th seesaw number $g_{i}(W)$ be defined as follows. The word $W_{i-1,i}$ can be uniquely written as $V_{1}\cdots V_{g_{i}(W)}$ for some $g_{i}(W)\geq 0$, such that $V_{\ell}$ is a non-empty braid word consisting only of $\sigma_{i-1}^{\pm 1}$ if $\ell+\epsilon$ is even, and only of $\sigma_{i}^{\pm 1}$ if $\ell+\epsilon$ is odd, for some fixed $\epsilon\in\{0,1\}$.
@@
(i). Let $C\subset S^{2}$ be a circle transverse to $D(W)$. After an isotopy pushing $C$ away from the crossings of $D(W)$, all intersection points $C\cap D(W)$ are on one of the strands. It follows that $C$ intersects each strand an even number of times. Thus if $C$ has two intersections with $D(W)$, then $C$ intersects only one strand (let us say the ii–th one), and does so twice. One checks that this is possible if and only if $i=1$, $i=n$, or $2\leq i\leq n-1$ and $g_{i}(W)\leq 3$. Moreover, $C$ is a decomposition circle if and only if $2\leq i\leq n-1$ and $g_{i}(W)\geq 2$.
(ii). Denote by $m_{-}$ and $m_{+}$ the number of times $W$ has been destabilized at the lower and upper end, respectively, to obtain $W^{\prime}$. Assume towards a contradiction that $g_{i}(W^{\prime})\leq 3$ for some $i$. Since $W^{\prime}$ is non-split, $g_{i}(W^{\prime})\geq 2$. Hence, as proven above, there is a decomposition circle $C^{\prime}$ of $D(W^{\prime})$ that intersects the ii–th strand of $W^{\prime}$ twice. This gives rise to a decomposition circle $C$ of $D(W)$ that intersects the $(i+m_{-})$–th strand of $W$ twice. This circle decomposes the link $\operatorname{cl}(W)$ as a connected sum of $L_{-}\coloneqq\operatorname{cl}(W_{1,i+m_{-}-1})$ and $L_{+}\coloneqq\operatorname{cl}(W_{i+m_{-},n})$. Since $W^{\prime}$ is non-split and cannot be destabilized, $W_{1,i+m_{-}-1}$ contains at least two copies of $\sigma_{m_{-}+1}^{\pm 1}$. Similarly, $\operatorname{cl}(W_{i+m_{-},n})$ contains at least two copies of $\sigma_{n-m_{+}}^{\pm 1}$. Using that homogeneous braids are visually trivial and visually split, it follows that $L_{+}$ and $L_{-}$ are non-trivial non-split links, and thus $C$ is an honest decomposition circle of $W$. ∎
@@ 5.3. Seifert surfaces of braids as Murasugi sums
Let us now describe the Seifert surface $\Sigma(W)$ for $\operatorname{cl}(\beta)$, obtained by applying Seifert’s algorithm to $D(W)$. The surface $\Sigma(W)$ consists of disks $D_{1},\dots,D_{n}$, such that $D_{i}$ has radius ni+1n-i+1, center $(0,0,i)$, is parallel to the xyxy–plane, and its boundary is the ii–th strand of $W$; as well as one twisted band for each crossing for each crossing of $W$. We call $D_{1}$ the bottom disk and $D_{n}$ the top disk. The following lemma gives a precise version of the well-known decomposition of $\Sigma(W)$ as Murasugi sum along the disks $D_{i}$. Examples can be seen in Figure 2, Figure 3, and Figure 4 in Section 2.
@@
Let us now explicitly define $P$. Recall from the definition of $g_{i}(W)\eqqcolon g$ that $W_{i-1,1}=V_{1}\cdots V_{g}$, where the words $V_{\ell}$ consist alternatingly only of $\sigma_{i-1}^{\pm 1}$, or only of $\sigma_{i}^{\pm 1}$. Now, let us pick points $p_{1},\ldots,p_{g}\in\partial\Sigma(W)$ that lie in this order on the circle $\partial D_{i}$, such that $p_{j}$ is between the spots where the twisted band corresponding to the last letter in $V_{j}$ and the twisted band corresponding to the first letter in $V_{j+1}$ are affixed to $D_{i}$ (where we see indexes modulo $g$ so that $V_{g+1}=V_{1}$). Let the polygon $P$ now be given a subset of the disk $D_{i}$, with vertices $p_{1},\ldots,p_{g-1}$ (and $p_{g}$ if $g$ is even), and edges given as chords of $D_{i}$. One checks that after an isotopy, the sides of $P$ indeed belong alternatingly to the boundary of the lower and the upper Murasugi summand of $\Sigma(W)$, thus proving (iii).
Finally, to show (iv), let us assume $g\geq 4$. Let $1\leq j\leq g$ and $j<g$ if $g$ is odd. Let us treat indexes modulo $g$ if $g$ is even, and modulo g1g-1 if $g$ is odd. Pick two twisted bands, such that both connect $D_{i}$ and $D_{i-1}$, or both connect $D_{i}$ and $D_{i+1}$: one of them affixed to $D_{i}$ between $p_{j}$ and $p_{j+1}$, the other between $p_{j+2}$ and $p_{j+3}$. Pick a simple closed curve $\gamma_{j}$ in $\Sigma(W)$ that consists of the core curves of those two twisted bands, and a chord in each of $D_{i}$ and $D_{i\pm 1}$. Observe that $\gamma_{j}$ is either contained in the lower or in the upper Murasugi summand of $\Sigma(W)$, depending on the sign in $D_{i\pm 1}$. Moreover, $\gamma_{j}$ intersects the side of $P$ between $p_{j}$ and $p_{j+1}$ exactly once. That is sufficient to prove that this side is essential. Overall, the sides of $P$ are alternatingly essential in the two Murasugi summands of $\Sigma(W)$; hence, the Murasugi sum is essential by definition. ∎
@@
Then $W$ satisfies the hypothesis of Lemma 5.1 (ii). So, for $W^{\prime}$ obtained from $W$ by destabilizing as often as possible, we have $g_{i}(W^{\prime})\geq 4$ for all $i$. Note that $\Sigma(W)$ and $\Sigma(W^{\prime})$ are isotopic. Hence it suffices to prove that $\Sigma(W^{\prime})$ is prime, i.e. that its associated open book for $S^{3}$ is prime. Also note that $W^{\prime}$ is itself homogeneous and non-split. Let $n$ be the number of strands of $W^{\prime}$.
An iterative application of Lemma 5.2 decomposes $\Sigma(W^{\prime})$ as an (n1)(n-1)–fold Murasugi sum with summands $\Sigma(W^{\prime}_{i,i})$. Because $W^{\prime}$ is homogeneous, every $W^{\prime}_{i,i}$ consists either of $k$ copies of $\sigma_{1}$, or of $k$ copies of $\sigma_{1}^{-1}$, with $k\geq 2$. So $\Sigma(W^{\prime}_{i,i})$ is the fiber surface of the $T(2,\pm k)$–torus link.
This Murasugi sum decomposition of $\Sigma(W^{\prime})$ is encoded by a tree of open books $T$ as defined in Section 3. Namely, the vertices of $T$ are the open books of the $\Sigma(W^{\prime}_{i,i})$; and there is an edge precisely between $\Sigma(W^{\prime}_{i-1,i-1})$ and $\Sigma(W^{\prime}_{i,i})$ for $2\leq i\leq n-1$. Note that the tree $T$ simply has the shape of a line. The edge between $\Sigma(W^{\prime}_{i-1,i-1})$ and $\Sigma(W^{\prime}_{i,i})$ is labeled by the summing region $P_{i}$ given by Lemma 5.2. Let us check that $T$ satisfies the hypothesis of a tree of open books. Firstly, the summing regions are disjoint, because $P_{i}$ is embedded in the ii–th disk $D_{i}$ of $\Sigma(W^{\prime})$. Secondly, since $g_{i}(W^{\prime})\geq 4$ for all $i$, it follows from Lemma 5.2 (iv) that $P_{i}$ is essential. So $T$ is indeed a tree of open books. The theorem now follows from Theorem 4.5. ∎